A rectangular garden is to be surrounded by a walkway of constant width. The garden's dimensions are 30 ft by 40 ft. The total area, garden plus walkway, is to be 1800 ft^2. What must be the width of the walkway to the nearest thousandth?
Found 2 solutions by checkley71, venugopalramana:
Answer by checkley71(8403)

(30+2X)(40+2X)=1800 OR 1200+80X+60X+4X^2=1800 OR 4X^2+140X-600=0 OR
X^2+35X-150=0 OR x=3.86 ft wide walkway (using the quadratic equation)
proof (30+2*3.86)(40+2*3.86)=1800 or (30+7.72)(40+7.72)=1800 or
37.72*47.72=1800 or 1800=1800

Answer by venugopalramana(3286)

A rectangular garden is to be surrounded by a walkway of constant width. The garden's dimensions are 30 ft by 40 ft. The total area, garden plus walkway, is to be 1800 ft^2. What must be the width of the walkway to the nearest thousandth?
WIDTH OF WALK WAY = X
LENGTH OF GARDEN WITH WALK WAY =40+X+X=40+2X
WIDTH..........................=30+X+X=30+2X
AREA =(2X+40)(2X+30)=1800
(X+20)(X+15)=450
X^2+35X+300-450=0
X^2+35X-150=0
X={-35+SQRT(35^2+4*150)}/2=3.860